• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name black

    Log entry time 09:18:30 on December 9,1998

    Entry number 7836

    keyword=beam current raised to 12 uA

    target temperatures ok. We are raising the beam current to 12 uA. MCC will
    do b-scope now while we run. Beam current was at 10 uA